solve for y in 5 + 6y = 8y - 21 and for p in 5p - 6 = 2(4-p)
Answer by Timnewman(323)

Hi dear,
The value of y=13 and p=2
HERE IS THE WORKING:
(1).5+6y=8y-21
colect the like term
8y-6y=5+21
2y=26
y=13.
soiving for p in 5p-6=2(4_p)
5p-6=8-2p
5p+2p=8+6
7p=14
p=2
Hope this helps?
Timnewman